OpenStudy (anonymous):

14. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept. Show your work.

OpenStudy (lynfran):

to find the x coordinate we let y=0 likewise, to find the y coordinate we let x=0

OpenStudy (lynfran):

so, y=-2(0)-3 and 0=-2x-3..just solve to get the respective coordinates

OpenStudy (lynfran):

y=-3 -2x-3=0, -2x=-3 x=-3/-2 x=3/2

OpenStudy (lynfran):

understand?

OpenStudy (anonymous):

Sorry i went to the bathroom ^ ^;

OpenStudy (lynfran):

ok...

OpenStudy (lynfran):

so what are the x intercept as in (x,y)... and the y intercept as in (x,y) coordinates

OpenStudy (anonymous):

1,2,3 for x and y is 5, 7, 9?

OpenStudy (lynfran):

no...please read my 3rd comment

OpenStudy (anonymous):

Alright.

OpenStudy (lynfran):

ok lets go over x.. intercept

OpenStudy (anonymous):

y = -2, 0, 3 x 0 = - 2x -3.

OpenStudy (lynfran):

when we want to find the x intercept of a function we let y=0 ie (x,0) im change the function and replacing it with y=5x-5 ok as and example so, 0=5x-5..we replace y and put 0 we now solve for x 5x-5=0 5x-5+5=+5 5x=5 we then divide both sides of the equation by 5 to solve for x 5x/5=5/5 x=1..

OpenStudy (anonymous):

Alright thanks. Do i need to make a graph?

OpenStudy (lynfran):

not always only if the question ask you to...so our x intercept for the given example would be (1,0)

OpenStudy (anonymous):

Alright. Thanks a lot you'r a self savor :)

OpenStudy (lynfran):

do u understand the y intecept part...if not i will go it over ok

OpenStudy (anonymous):

It intercepts at 0 right?

OpenStudy (lynfran):

we let the x intercept =0 ie (0,y) and we solve for y...what would be the y intercept for the given example i just give?

OpenStudy (lynfran):

this one y=5x-5

OpenStudy (anonymous):

x is equal to 0?

OpenStudy (lynfran):

yes we let x=o and solve for y y=5(0)-5

OpenStudy (anonymous):

Wouldn't that be 0 too?

OpenStudy (lynfran):

no

OpenStudy (lynfran):

y=5*0-5 y=0-5 y=-5

OpenStudy (anonymous):

Okay. Think i got it now 5*0 = 0 and 0 subtracted by -5 is just negative 5.

OpenStudy (anonymous):

-5*

OpenStudy (lynfran):

yes now put it in coordinate form..(x,y) we get..

OpenStudy (anonymous):

x= 0 and y= -5

OpenStudy (lynfran):

very good (0,-5)

OpenStudy (anonymous):

Thanks. I suck at math. XD

OpenStudy (lynfran):

no problem but remember that was the intercepts for the example i give not for the problem you give ok

OpenStudy (anonymous):

Yeah totally understand. That's why i try in solve them before i post them unless i really have no clue where to start.

OpenStudy (lynfran):

ok just to note also in the case of sketching a graph..the x intercept is where the line/curve cuts the x axis on the graph...and the y intercept is where the line/curve cuts the y axis on the graph.

OpenStudy (anonymous):

Alright kool.